OpenStudy (anonymous):

Choose the value of the y determinant (Dy) in the following system. 2x + 3y = 4 5x – y = 8

OpenStudy (anonymous):

–17 –28 –13 –4

OpenStudy (nenadmatematika):

-4

razor99 (razor99):

thats simultaneous equation

OpenStudy (anonymous):

how did you find that?

OpenStudy (nenadmatematika):

when you calculate the y determinant Dy, first column is consisting of coefficients at x, and second column is consisting of elements on the right side of equality sign. So it is: \[Dy=\left[\begin{matrix}2 & 4 \\ 5 &8\end{matrix}\right]\] =2*8-4*5=16-20=-4

OpenStudy (anonymous):

thanks so much!

OpenStudy (nenadmatematika):

you're welcome carlyn :D